Circular No. 6452

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

1996 PW MPEC 1996-P03 contains observations and a preliminary orbit for an

asteroidal object with a record high e. 1996 PW was identified and recognized as unusual by Gareth Williams, Minor Planet Center, Harvard-Smithsonian Center for Astrophysics, in two nights (Aug. 9 and 12) of general asteroidal data from the NEAT program, carried out under the direction of Eleanor Helin, Jet Propulsion Laboratory, using images obtained with the GEODSS telescope at Haleakala. More recent observations demonstrate that e > 0.99, with other elements T = 1996 Aug. 8, ω = 182^{o}, Ω = 144^{o}, i = 30^{o} q = 2.537 AU, H = 14.0. Imaging by D. Jewitt on Aug. 16 with the University of Hawaii’s 2.2-m reflector showed the object as a point source, devoid of coma in integrations that would easily show the coma of (2060) = comet 95P/Chiron, for example. Photometry by D. Rabinowitz, Jet Propulsion Laboratory, at the Table Mountain Observatory on Aug. 14.4 UT yields V = 17.9, V − R = +0.31 ± 0.05, V − I = +0.60 ± 0.05 (assuming solar values V − R = +0.367, V − I = +0.705).

CYGNUS X-2 A. P. Smale, NASA/Goddard Space Flight Center and Universities

Space Research Association; E. Kuulkers, Astrophysics Division, ESTEC, European Space Agency; and R. A. D. Wijnands, Astronomical Institute ‘Anton Pannekoek’, University of Amsterdam, report: “The public RXTE ASM lightcurve (covering the interval 1996 Feb. 22–Aug. 8) of the 9.8-day low-mass x-ray-binary Cyg X-2 shows a pronounced intensity modulation with a period of 77.0±1.0 days and full amplitude of 40 percent. A shallower (20 percent) secondary minimum is also observed. The spectral hardness and source intensity are anticorrelated on this cycle. We find that this longterm periodicity can be detected independently in the HEASARC archival datasets from Vela 5B (see also Smale and Lochner 1992, Ap.J. 395, 582) and the Ariel V ASM. From these three datasets we derive a single overall ephemeris from 1969 to 1996 of HJD 2442202.5 (±5.0) + 77.13 (±0.05)E (where phase zero is the primary minimum). The occurrence of different intensity states in the Einstein, EXOSAT and Ginga data from Cyg X-2 (Kuulkers et al. 1996, A.Ap. 311, 197) is also consistent with this 77.1-day periodicity, which we tentatively associate with the precession of a tilted accretion disk in the system.”

1996 August 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden